OpenStudy (anonymous):

Does anyone know the answer to this question? 2√6+3 √96

OpenStudy (anonymous):

not trying to be funny, but are you not allowed to use a calculator?

OpenStudy (anonymous):

Yes, but I don't know how to type it in. I don't know where to start really :P

jhonyy9 (jhonyy9):

2sqrt6 +3sqrt(16*6) =2sqrt6 +3*4sqrt6 =2sqrt6 +12sqrt6 =14sqrt6

OpenStudy (anonymous):

helps to know \(96=16\times 6\) so \(\sqrt{96}=4\sqrt{6}\) then you can combine like terms

OpenStudy (anonymous):

Well that was one of my choices, so I applaud you :D Thank you guys !

OpenStudy (anonymous):

\[2\sqrt{6}+12\sqrt{6}=....\]

OpenStudy (anonymous):

I went through the steps and it sorta makes sense, but I am horrible at math -.-

OpenStudy (anonymous):

no panic.when using your calculator use brackets a lot to avoid problems. write it like this: \[(2\times \sqrt{6}) + (3\times \sqrt{96})=\] press the squarre root button before you type the digits you want to put under it